OpenStudy (anonymous):

if k=1/2 mvsquare solve for V

OpenStudy (shinalcantara):

\[k = \frac{ 1 }{ 2 } m v^2\] rearranging you'll have \[v^2 = \frac{ 2k }{ m }\] \[v = \sqrt{\frac{ 2k }{ m }}\]